# Kaidi Zhang1, Yunmiao Zhang2, Honghai Yu2, Yan Hu1, Benyou Wang122footnotemark: 2

### v4

@@ Appendix H Prompt
{user_profile[‘prompt’} You are a {user_strategy} investor(fundamental or technical)
@@
Identity-Enhancing Prompt Identity-Enhancing Prompt:
@@
Current date is: {format_date(cur_date)}, is ({“Trading Day” if is_trading_day else “Non-Trading Day”}). Your previous day’s belief is: {belief}
@@
Current total assets: {user_profile[“total_value”] / 1000:,.2f} thousand yuan Available cash: {user_profile[“current_cash”] / 1000:,.2f} thousand yuan Cumulative rate of return: {user_profile[“return_rate”]}%
@@
Forum Checking Prompt Forum Checking Prompt:
@@
Now you are browsing a forum, and you need to make a decision for each post, deciding whether to perform an action on the post. Your decision should be in line with your investment style and persona. Here is the information of the current post: {post_id} {post_content} The post quotes the following content: {root_content} Please decide whether to perform an action on this post based on the above information. You can choose one of the following actions: Repost: You think this post is worth sharing with more people, and you can add your comments Unlike: You think this post is not worth liking Like: You think this is a valuable post Please note that your analysis should be based on the posts you see. Please output your decision in the following format: <action >Operation type </ action >   <reason >Output your reason </ reason >
@@
Important News Analysis Prompt Important News Analysis Prompt:
I will provide you with filtered news that is highly time-sensitive and important. This news is **public news**. Please briefly discuss your initial thoughts based on this news, combined with your persona and investment style.
News List: {formatted_news}
@@
Initial News Query Prompt Initial News Query Prompt:
Based on historical trading and system recommendations, all the assets and corresponding industries you are currently paying attention to are as follows: {stock_details}(including the stocks we recommend to users through our recommendation system)
@@
Based on your investment preferences and the current market situation, please consider the following questions: What type of information do you hope to obtain from the news? (e.g., market trends, industry information, etc.) Do you have specific keywords or topics that you need to further understand?
@@
News Query Formulation Prompt News Query Formulation Prompt:
Based on the questions you just summarized, you now need to input the content you want to query to retrieve relevant news and announcements. Your output should be in YAML format:
queries: # list[str],required, each string represents an independent query question, sorted by importance, the question should be a specific question about a stock or industry, for example, it should be <White wine consumption trend>, not <White wine consumption upgrade> - Your question 1 - Your question 2 stock_id: # list[str], optional, each string represents the stock code of a company you want to query - Stock code 1 - Stock code 2
@@
Belief Update Prompt Belief Update Prompt:
@@
{old_belief} Based on the news and announcements you searched for, the posts you browsed, combined with your persona and previous belief, please describe your new belief in the first person in one paragraph. Please output a paragraph directly, without any additional structure or headings. Your answer should include the following: Market Trend: Please describe your view on the general direction of the market in the next month at the current time. Market Valuation: Please describe your view on the current market valuation at the current time. Economic Condition: Please describe your view on the future macroeconomic trends at the current time. Market Sentiment: Please describe your view on the current market sentiment at the current time. Self-Evaluation: Combining your historical trading performance and investment style at the current time, please describe your evaluation of your self-investment level. Please try to make your answer natural and fluent, and avoid mechanical template-based expressions. Please output plain text format directly.
@@
Potential Index Selection Prompt Potential Index Selection Prompt:
@@
Your Holding Status: {current_stock_details} Industries You Are Currently Following: {‘, ’.join(fol_ind)} System-Recommended Industry indexes: {potential_stock_details} Your Current Belief: {belief if belief else “None”}
Please return the list of indexes you want to focus on today and the reasons in YAML format: Please note: The reason field should include your reasons for choosing these indexes, such as based on your persona and investment style, or based on your belief, explained in a paragraph selected_index: # Select all indexes you potentially want to trade, just output the index codes (English codes), do not output index names - Index code 1 - Index code 2 reason:
@@
Querying Stock Data Prompt Querying Stock Data Prompt:
@@
Summary Market Quotes for Relevant indexes on the Previous Trading Day are as follows: {stock_summary} Your Holding Information for Relevant indexes is as follows: {positions_info} Based on your role and the available data, determine whether additional data is needed for analysis. If needed, you can directly query relevant data to assist in decision-making.
Note: Please naturally obtain data indicators that you consider valuable during the analysis process. Data acquisition is part of the analysis, but please also be efficient and select the most relevant indicators to support your analysis, otherwise you will be penalized. Please pay attention to your persona and investment style. The analysis and reasoning process should start from the persona and remain natural and coherent. As a reminder: You are a {user_strategy investor Your current portfolio’s total return rate is {return_rate}%, your current total assets are {total_value:,} RMB, and your current cash balance is {current_cash:,} RMB. The reason field should include your reasons for choosing these indicators, such as based on your persona and investment style, or based on your belief, explained in a paragraph.
@@
Output in the following YAML format: indicators: # List(strs), indicating the indicators you think need to be filtered - indicator1 - indicator 2 start_date: ’%Y-%m-%d # Start time for querying end_date: ’%Y-%m-%d # End time for querying reason:
@@
Trading Decision Making Prompt Trading Decision Making Prompt(Step 1):
Now it’s time to make the final trading decision. Please base your analysis on all the information you have obtained previously, combined with your investment style and persona. First, conduct an analysis, and then make specific trading decisions and provide your reasons for each industry index. A: Please analyze the industry indices you want to trade: {‘, ’.join(stocks_to_deal)} based on all the information you have obtained previously, including news, announcements, market data, industry data, and additional stock data.
Please Include the Following: 1. Overall Market Analysis: Views on the current overall market situation. Main trends and possible changes. 2. Impact of News and Announcements: Analysis of the impact of important news and announcements on the market and individual stocks. Are there any major events that may change market trends? 3. Industry Index Analysis: Detailed analysis for each industry index of interest: Current performance and technical analysis. Fundamentals and future expectations. Buy, hold, or sell recommendations. 4. Risk Assessment: Main risk points in the current market. Risk assessment for each asset. Note: Please conduct the analysis in conjunction with the set persona and investment style. Please conduct the analysis in conjunction with all the information you have obtained previously. The output content should be clear, concise, and easy to understand. Output Format for Analysis: Please output the analysis results in natural language to ensure clear logic and complete structure.
@@
Trading Decision Making Prompt Trading Decision Making Prompt(Step 2):
@@
Trading Related Information: Remaining available position (relative to total assets): {available_position:.2f}Specific Information for Each Industry Index: {chr(10).join(stock_info)} Trading Rules: 1. The transaction price must be between the lower and upper limit prices. 2. Please fill in the trading position and price. If you choose to hold, the position for preparation of the transaction is 0.
Precautions: Important: All decisions must be consistent with your investment style and persona, but you should also make appropriate adjustments based on the actual situation of the day. Ensure that each trading decision is within the price and position limits. If you choose to hold, trading_position should be equal to 0. If the current position of an industry index is 0, it means it is a system-recommended industry index, and you can only choose to buy or hold. If you choose to sell or buy, please pay attention to the setting of your trading_position (if you choose to sell, then trading_position cannot exceed the current position). If the setting is unreasonable, it may lead to transaction failure. You need to choose the target_price according to your expectations, instead of setting it to yesterday’s closing price. Important: trading_position represents the percentage of your total assets that you plan to use for this transaction, and it is always positive.
@@
Social Media Posting Prompt Social Media Posting Prompt:
@@
1. Post Content: - Your post must fall into one of the following three categories: - type1: Commentary on an Event - Cite specific news or announcements, express your opinions and analysis, and integrate personal insights with a conversational tone. - type2: Summary of Your Recent Trading Behavior - Summarize your latest trading actions, explain the underlying logic and outcomes in detail, and share your emotions and reflections. - type3: Market Outlook - Based on current market information, predict future trends or investment opportunities while expressing your expectations or concerns about the future. 2. Content Requirements: - Align your post with your investment style and persona, incorporating your trading decisions, perspectives, and analysis. Choose the most suitable post type and writing style. - Keep the post between 100-200 words. - Clearly indicate the post type (type1/type2/type3). 3. Belief Summary: - Your prior belief was: {old_belief}
@@
- Market Trend - Describe your outlook on market direction for the next month. - Market Valuation - Express your opinion on the current market valuation. - Economic Conditions - Share your expectations regarding macroeconomic trends. - Market Sentiment - Describe your perception of the current market sentiment. - Self-Assessment - Reflect on your historical trading performance and investment style to evaluate your investment ability. 4. Output Format: - The output should follow the YAML format below: post: Your post content type: type1/type2/type3 # Post type, required, string format belief: Your Belief Summary

### v5

@@ Appendix F Prompt
  • {user_profile[‘prompt’]}
  • You are a {user_strategy} investor (fundamental or technical)
@@
{promptbox}
Identity-Enhancing Prompt for TwinMarket Agents Identity-Enhancing Prompt:
@@
  • Current date is: {format_date(cur_date)}, is ({“Trading Day” if is_trading_day else “Non-Trading Day”}).
  • Your previous day’s belief is: {belief}
@@
  • Current total assets: {user_profile[“total_value”] / 1000:,.2f} thousand yuan
  • Available cash: {user_profile[“current_cash”] / 1000:,.2f} thousand yuan
  • Cumulative rate of return: {user_profile[“return_rate”]}%
@@
{promptbox}
Forum Checking Prompt for TwinMarket Agents Forum Checking Prompt:
@@
Now you are browsing a forum, and you need to make a decision for each post, deciding whether to perform an action on the post. Your decision should be in line with your investment style and persona.
Here is the information of the current post:
  • {post_id}
  • {post_content}
The post quotes the following content: {root_content}
Please decide whether to perform an action on this post based on the above information.
You can choose one of the following actions:
  • Repost: You think this post is worth sharing with more people, and you can add your comments
  • Unlike: You think this post is not worth liking
  • Like: You think this is a valuable post
Please note that your analysis should be based on the posts you see.
Please output your decision in the following format:
  • <action> Operation type </action> <reason> Output your reason </reason>
@@
{promptbox}
Important News Analysis Prompt for TwinMarket Agents Important News Analysis Prompt:
I will provide you with filtered news that is highly time-sensitive and important. This news is public news. Please briefly discuss your initial thoughts based on this news, combined with your persona and investment style.
News List:
  • {formatted_news}
@@
{promptbox}
Initial News Query Prompt for TwinMarket Agents Initial News Query Prompt:
Based on historical trading and system recommendations, all the assets and corresponding industries you are currently paying attention to are as follows:
{stock_details} (including the stocks we recommend to users through our recommendation system)
@@
Based on your investment preferences and the current market situation, please consider the following questions:
  • What type of information do you hope to obtain from the news? (e.g., market trends, industry information, etc.)
  • Do you have specific keywords or topics that you need to further understand?
@@
{promptbox}
News Query Formulation Prompt for TwinMarket Agents News Query Formulation Prompt:
Based on the questions you just summarized, you now need to input the content you want to query to retrieve relevant news and announcements.
Your output should be in YAML format:
queries: # list[str],required, each string represents an independent query question, sorted by importance, the question should be a specific question about a stock or industry, for example, it should be <White wine consumption trend>, not <White wine consumption upgrade>
- Your question 1
- Your question 2
stock_id: # list[str], optional, each string represents the stock code of a company you want to query
- Stock code 1
- Stock code 2
@@
{promptbox}
Belief Update Prompt for TwinMarket Agents Belief Update Prompt:
@@
  • {old_belief}
Based on the news and announcements you searched for, the posts you browsed, combined with your persona and previous belief, please describe your new belief in the first person in one paragraph. Please output a paragraph directly, without any additional structure or headings. Your answer should include the following:
  • Market Trend: Please describe your view on the general direction of the market in the next month at the current time.
  • Market Valuation: Please describe your view on the current market valuation at the current time.
  • Economic Condition: Please describe your view on the future macroeconomic trends at the current time.
  • Market Sentiment: Please describe your view on the current market sentiment at the current time.
  • Self-Evaluation: Combining your historical trading performance and investment style at the current time, please describe your evaluation of your self-investment level.
Please try to make your answer natural and fluent, and avoid mechanical template-based expressions. Please output plain text format directly.
@@
{promptbox}
Potential Index Selection Prompt for TwinMarket Agents Potential Index Selection Prompt:
@@
Your Holding Status:
  • {current_stock_details}
Industries You Are Currently Following:
  • {‘, ’.join(fol_ind)}
System-Recommended Industry indexes:
  • {potential_stock_details}
Your Current Belief:
  • {belief if belief else “None”}
Please return the list of indexes you want to focus on today and the reasons in YAML format:
  • Please note: The reason field should include your reasons for choosing these indexes, such as based on your persona and investment style, or based on your belief, explained in a paragraph
selected_index: # Select all indexes you potentially want to trade, just output the index codes (English codes), do not output index names
- Index code 1
- Index code 2
reason:
@@
{promptbox}
Querying Stock Data Prompt for TwinMarket Agents Querying Stock Data Prompt:
@@
Summary Market Quotes for Relevant indexes on the Previous Trading Day are as follows:
  • {stock_summary}
Your Holding Information for Relevant indexes is as follows:
  • {positions_info}
Based on your role and the available data, determine whether additional data is needed for analysis. If needed, you can directly query relevant data to assist in decision-making.
Note:
  • Please naturally obtain data indicators that you consider valuable during the analysis process. Data acquisition is part of the analysis, but please also be efficient and select the most relevant indicators to support your analysis, otherwise you will be penalized.
  • Please pay attention to your persona and investment style. The analysis and reasoning process should start from the persona and remain natural and coherent.
  • As a reminder: You are a {user_strategy} investor
  • Your current portfolio’s total return rate is {return_rate}%, your current total assets are {total_value:,} RMB, and your current cash balance is {current_cash:,} RMB.
  • The reason field should include your reasons for choosing these indicators, such as based on your persona and investment style, or based on your belief, explained in a paragraph.
@@
Output in the following YAML format:
indicators: # List(strs), indicating the indicators you think need to be filtered
- indicator1
- indicator 2
start_date: ’%Y-%m-%d # Start time for querying
end_date: ’%Y-%m-%d # End time for querying
reason:
@@
{promptbox}
Trading Decision Making Prompt (Step 1) for TwinMarket Agents Trading Decision Making Prompt (Step 1):
Now it’s time to make the final trading decision. Please base your analysis on all the information you have obtained previously, combined with your investment style and persona. First, conduct an analysis, and then make specific trading decisions and provide your reasons for each industry index. Please analyze the industry indices you want to trade: {‘, ’.join(stocks_to_deal)} based on all the information you have obtained previously, including news, announcements, market data, industry data, and additional stock data.
Please Include the Following:
  1. 1.
    Overall Market Analysis:
    • Views on the current overall market situation.
    • Main trends and possible changes.
  2. 2.
    Impact of News and Announcements:
    • Analysis of the impact of important news and announcements on the market and individual stocks.
    • Are there any major events that may change market trends?
  3. 3.
    Industry Index Analysis:
    • Detailed analysis for each industry index of interest:
      • Current performance and technical analysis.
      • Fundamentals and future expectations.
      • Buy, hold, or sell recommendations.
  4. 4.
    Risk Assessment:
    • Main risk points in the current market.
    • Risk assessment for each asset.
Note:
  • Please conduct the analysis in conjunction with the set persona and investment style.
  • Please conduct the analysis in conjunction with all the information you have obtained previously.
  • The output content should be clear, concise, and easy to understand.
Output Format for Analysis: Please output the analysis results in natural language to ensure clear logic and complete structure.
@@
{promptbox}
Trading Decision Making Prompt (Step 2) for TwinMarket Agents Trading Decision Making Prompt (Step 2):
@@
Trading Related Information:
  • Remaining available position (relative to total assets): {available_position:.2f}%
Specific Information for Each Industry Index:
  • {chr(10).join(stock_info)}
Trading Rules:
  1. 1.
    The transaction price must be between the lower and upper limit prices.
  2. 2.
    Please fill in the trading position and price. If you choose to hold, the position for preparation of the transaction is 0.
Precautions:
  • Important: All decisions must be consistent with your investment style and persona, but you should also make appropriate adjustments based on the actual situation of the day.
  • Ensure that each trading decision is within the price and position limits.
  • If you choose to hold, trading_position should be equal to 0.
  • If the current position of an industry index is 0, it means it is a system-recommended industry index, and you can only choose to buy or hold.
  • If you choose to sell or buy, please pay attention to the setting of your trading_position (if you choose to sell, then trading_position cannot exceed the current position). If the setting is unreasonable, it may lead to transaction failure. You need to choose the target_price according to your expectations, instead of setting it to yesterday’s closing price.
  • Important: trading_position represents the percentage of your total assets that you plan to use for this transaction, and it is always positive.
@@
{promptbox}
Social Media Posting Prompt for TwinMarket Agents Social Media Posting Prompt:
@@
1. Post Content:
  • Your post must fall into one of the following three categories:
    • type1: Commentary on an Event - Cite specific news or announcements, express your opinions and analysis, and integrate personal insights with a conversational tone.
    • type2: Summary of Your Recent Trading Behavior - Summarize your latest trading actions, explain the underlying logic and outcomes in detail, and share your emotions and reflections.
    • type3: Market Outlook - Based on current market information, predict future trends or investment opportunities while expressing your expectations or concerns about the future.
2. Content Requirements:
  • Align your post with your investment style and persona, incorporating your trading decisions, perspectives, and analysis. Choose the most suitable post type and writing style.
  • Keep the post between 100-200 words.
  • Clearly indicate the post type (type1/type2/type3).
3. Belief Summary:
  • Your prior belief was: {old_belief}
@@ 3. Belief Summary:
    • Market Trend - Describe your outlook on market direction for the next month.
    • Market Valuation - Express your opinion on the current market valuation.
    • Economic Conditions - Share your expectations regarding macroeconomic trends.
    • Market Sentiment - Describe your perception of the current market sentiment.
    • Self-Assessment - Reflect on your historical trading performance and investment style to evaluate your investment ability.
4. Output Format:
  • The output should follow the YAML format below:
post: Your post content
type: type1/type2/type3 # Post type, required, string format
belief: Your Belief Summary